# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ IV-B2 CRLB of CA-enabled MIMO-OFDM system
The CRLBs of range and velocity estimations of the proposed ISAC system are derived. For simplicity, two CCs occupy the same number of resource elements (REs), i.e., $N_{1}=N_{2}=N$ and $M_{1}=M_{2}=M$.
According to (12), the CRLBs of $\tau_{0}$ and $\gamma$ for the proposed CA-enabled MIMO-OFDM ISAC system are expressed as (56) and (57), respectively.
$CRLB_{\text{CA}}(\tau_{0})=\frac{\Gamma_{\text{ca}}N\left[MN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{m}-\left(\Theta_{B}^{p,m}\right)^{2}\right]}{MN_{\text{U}}},$ (56)
$CRLB_{\text{CA}}(\gamma)=\frac{\Gamma_{\text{ca}}M\left[NN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{n}-\left(\Theta_{B}^{p,n}\right)^{2}\right]}{NN_{\text{U}}},$ (57)
$\Gamma_{\text{ca}}=\frac{1}{\varpi\cdot\left(2\pi\right)^{2}}\frac{1}{MNN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{n}\Theta_{B}^{m}+2\Theta_{B}^{p,n}\Theta_{B}^{p,m}\Theta_{B}^{m,n}-N\left(\Theta_{B}^{p,m}\right)^{2}\Theta_{B}^{n}-M\left(\Theta_{B}^{p,n}\right)^{2}\Theta_{B}^{m}-N_{\text{U}}\left(\Theta_{B}^{m,n}\right)^{2}\Theta_{B}^{p}},$ (58)
where $\Gamma_{\text{ca}}$ is expressed in (58), $\Theta_{B}^{n}=\sum\limits_{n}\left(n\left(\Delta f^{1}+\Delta f^{2}\right)\right)^{2}$, $\Theta_{B}^{m}=\sum\limits_{m}\left(m\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right)\right)^{2}$, $\Theta_{B}^{p}=\sum\limits_{p}\left(p\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\right)^{2}$, and
$\Theta_{B}^{m,n}=\sum\limits_{m}\sum\limits_{n}mn\left(\Delta f^{1}+\Delta f^{2}\right)\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right),$
@@
$\Theta_{B}^{p,n}=\sum\limits_{p}\sum\limits_{n}pn\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\left(\Delta f^{1}+\Delta f^{2}\right),$
@@
$\Theta_{B}^{p,m}=\sum\limits_{p}\sum\limits_{m}pm\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right).$
Substituting $\Delta f^{2}=Q\Delta f^{1}$ and $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ mentioned in Section III-B into (56) and (57), the simplest close-form expressions for CRLBs of $\tau_{0}$ and $\gamma$ are
$CRLB_{\text{CA}}(\tau_{0})=\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2M-1)}{36}-\frac{(N_{\text{U}}-1)(M-1)}{16}\right]}{\left(\Delta f^{1}\right)^{2}\left(Q+1\right)^{2}\left(N-1\right)},$ (59)
$CRLB_{\text{CA}}(\gamma)=\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2N-1)}{36}-\frac{(N_{\text{U}}-1)(N-1)}{16}\right]}{\left(2f_{\mathrm{C}}^{1}T^{1}\right)^{2}\left(M-1\right)},$ (60)
where $\Gamma_{\text{CA}}$ is expressed in (61).
$\Gamma_{\text{CA}}=\frac{1}{\varpi(2\pi)^{2}MNN_{\text{U}}}\times$ (61)
$\frac{1}{\frac{(2N_{\text{U}}-1)}{6}\frac{(2N-1)}{6}\frac{(2M-1)}{6}+(N_{\text{U}}-1)\left[\frac{(N-1)(M-1)}{32}-\frac{(M-1)(2N-1)}{96}-\frac{(N-1)(2M-1)}{96}\right]-\left(\frac{(M-1)(N-1)}{16}\right)\left(\frac{(2N_{\text{U}}-1)}{6}\right)}.$
According to $\tau_{0}=\frac{2R}{c}$ and $\gamma=\frac{2v_{0}}{c}$, the CRLBs of $R$ and $v_{0}$ of CA-enabled MIMO-OFDM ISAC system are
$CRLB_{\text{CA}}(R)=\frac{c^{2}}{4}\cdot\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2M-1)}{36}-\frac{(N_{\text{U}}-1)(M-1)}{16}\right]}{\left(\Delta f^{1}\right)^{2}\left(Q+1\right)^{2}\left(N-1\right)},$ (62)
$CRLB_{\text{CA}}(v_{0})=\frac{c^{2}}{4}\cdot\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2N-1)}{36}-\frac{(N_{\text{U}}-1)(N-1)}{16}\right]}{\left(2f_{\mathrm{C}}^{1}T^{1}\right)^{2}\left(M-1\right)}.$ (63)
@@ V Simulation Results and Analysis
In this section, the simulation results are provided to verify the high-spend communication and high-accuracy sensing capabilities of the proposed CA-enabled MIMO-OFDM ISAC system. Simulation parameters are detailed in Table II. The simulation results are obtained with 1000 times Monte Carlo simulations.
TABLE II: Simulation parameters [7, 29, 30].
@@
$f_{\mathrm{C}}^{1}$ Carrier frequency of low-frequency band 5.9 GHz [7] $f_{\mathrm{C}}^{2}$ Carrier frequency of high-frequency band 24 GHz [7]
$M_{1}$ Number of OFDM symbols on low-frequency band 64 $M_{2}$ Number of OFDM symbols on high-frequency band 64
$N_{1}$ Number of subcarriers on low-frequency band 512 $N_{2}$ Number of subcarriers on high-frequency band 512
$T^{1}$ Total symbol duration on low-frequency band 39.5 $\mathrm{\mu}$s $T^{2}$ Total symbol duration on high-frequency band 9.7 $\mathrm{\mu}$s
$\Delta f^{1}$ Subcarrier spacing on low-frequency band 30 kHz [30] $\Delta f^{2}$ Subcarrier spacing on high-frequency band 120 kHz [30]
$N_{\text{T}}$ Number of transmit antenna of BS 4 $N_{\text{R}}$ Number of received antenna of BS 4
$N_{\text{U}}$ Number of received antenna of UE {4, 5, 6} $\theta_{\text{Rx}},\theta_{\text{Tx}}$ AoA and AoD of target $30^{\circ}$
$r_{0}$ Range of target 117 m $v_{0}$ Velocity of target 13 m/s
@@ V-A Performance of Communication
According to (43) and (46) derived in Section IV-A, the communication MIs of CA-enabled MIMO-OFDM ISAC system and MIMO-OFDM ISAC system under various SNR and $N_{\text{U}}$ are simulated in Fig.  5.
@@
In Fig. 5, the communication MI of the CA-enabled MIMO-OFDM ISAC signal is larger than that of the MIMO-OFDM ISAC signal under different $N_{\text{U}}$.
@@ V-B Performance of Sensing
In this section, the CRLBs and RMSEs of range and velocity estimations are simulated to demonstrate the high-accuracy sensing performance of the proposed CA-enabled MIMO-OFDM ISAC system.
V-B1 CRLB
@@
  • Upon observing Fig. 7(b), the high-frequency MIMO-OFDM ISAC signal and low-frequency MIMO-OFDM ISAC signal are overlapped, which also exist in the OFDM ISAC signals. The reason is as follows. According to (63), for the ISAC signals occupying the same number of REs and satisfying $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ in simulation parameters, the CRLBs of velocity estimation are identical.
@@
V-B2 RMSE
The data-level and signal-level fusion methods are compared in Fig. 8, where the the following phenomena are revealed.
  • In Fig. 8, the RMSE of symbol-level fusion method is consistently lower than that of the data-level fusion method, indicating the superiority of the proposed symbol-level fusion method over the data-level fusion method.
  • Compared with the data-level fusion method, the accuracy of range and velocity estimations under symbol-level fusion method is improved by up to 66.91% and 95.67%, respectively.
Then, the RMSEs for range and velocity estimations of the proposed CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are simulated in Fig. 9. There are several assumptions.
@@ V-B2 RMSE
    The theoretical RMSEs of range and velocity estimations are expressed in (64) and (65), respectively [31].
@@
(a) RMSE for range estimation
(b) RMSE for velocity estimation
@@
Fig. 9(a) reveals the following phenomena.
  • With the increase of SNR, the RMSEs of CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are stable, because the accuracy of sensing method is limited by the grid size.
  • Comparing with the RMSEs of MIMO-OFDM ISAC signals, the RMSE of CA-enabled MIMO-OFDM ISAC signal is lower, indicating the superior performance of range estimation of the proposed CA-enabled MIMO-OFDM ISAC signal over the MIMO-OFDM ISAC signal.
  • When SNR is smaller than $-20$ dB, comparing with the RMSE of low-frequency MIMO-OFDM ISAC signal, the RMSE of high-frequency MIMO-OFDM ISAC signal is higher. In the simulation, we employed a typical path loss model and set the received SNR of high-frequency signal 5 dB lower than that of low-frequency signal [32]. Therefore, for comparison under the same received SNR, the curve of high-frequency MIMO-OFDM ISAC signal needs to shift 5 dB to the left.
According to (65), since $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ in the simulation, the differences in the RMSEs of high and low frequency signals are only the received SNR.
(a) RMSE for range estimation
(b) RMSE for velocity estimation

### v2

@@ IV-B2 CRLB of CA-enabled MIMO-OFDM system
The CRLBs of parameter estimations of the proposed ISAC system are derived.
Before deriving, we need to know that for independent observed data, the total likelihood function is the product of the likelihood functions of each independent observation, and the total log-likelihood function is the sum of the log-likelihood functions of each independent observation. Therefore, for two independent observations, the total Fisher Information Matrix (FIM) is the sum of the FIMs of the two independent observations [44].
Upon observing (12), the received echo signals on high and low-frequency bands are independent. Therefore, by linearly summing the respective FIM of the received echo signals on high and low-frequency bands, the CRLB for the CA-enabled MIMO-OFDM system can be derived.
The FIM of the $b$-th CC is expressed as
$\mathbf{F}_{b}=\begin{bmatrix}F_{\sin\theta_{i,\text{Rx}}}^{b}&F_{\sin\theta_{i,\text{Rx}},\tau_{i,0}}^{b}&F_{\sin\theta_{i,\text{Rx}},\gamma_{i}}^{b}\\ F_{\tau_{i,0},\sin\theta_{i,\text{Rx}}}^{b}&F_{\tau_{i,0}}^{b}&F_{\tau_{i,0},\gamma_{i}}^{b}\\ F_{\gamma_{i},\sin\theta_{\text{Rx}}}^{b}&F_{\gamma_{i},\tau_{i,0}}^{b}&F_{\gamma_{i}}^{b}\end{bmatrix},$ (51)
where $F_{\sin\theta_{i,\text{Rx}}}^{b}$, $F_{\tau_{i,0}}^{b}$, and $F_{\gamma_{i}}^{b}$ are expressed in (52), (53), and (54), respectively; $F_{\sin\theta_{i,\text{Rx}},\tau_{i,0}}^{b}=F_{\tau_{i,0},\sin\theta_{i,\text{Rx}}}^{b}$ is expressed in (55); $F_{\sin\theta_{i,\text{Rx}},\gamma_{i}}^{b}=F_{\gamma_{i},\sin\theta_{i,\text{Rx}}}^{b}$ is expressed in (56); $F_{\tau_{i,0},\gamma_{i}}^{b}=F_{\gamma_{i},\tau_{i,0}}^{b}$ is expressed in (57).
$F_{\sin\theta_{i,\text{Rx}}}^{b}=\frac{\left|A\right|^{2}}{\sigma^{2}}\sum_{p}^{N_{\text{R}}}\sum_{m}^{M_{b}}\sum_{n}^{N_{b}}\left(2\pi p\frac{d_{\text{r}}}{\lambda^{b}}\right)^{2}=\frac{4\pi^{2}\left|A\right|^{2}}{\sigma^{2}}\left(\frac{d_{\text{r}}}{\lambda^{b}}\right)^{2}\frac{\left(2N_{\text{R}}-1\right)(N_{\text{R}}-1)N_{\text{R}}N_{b}M_{b}}{6},$ (52)
$F_{\tau_{i,0}}^{b}=\frac{\left|A\right|^{2}}{\sigma^{2}}\sum_{p}^{N_{\text{R}}}\sum_{m}^{M_{b}}\sum_{n}^{N_{b}}\left(2\pi n\Delta f^{b}\right)^{2}=\frac{4\pi^{2}\left|A\right|^{2}\left(\Delta f^{b}\right)^{2}}{\sigma^{2}}\frac{\left(2N_{b}-1\right)\left(N_{b}-1\right)N_{b}N_{\text{R}}M_{b}}{6},$ (53)
$F_{{\gamma_{i}}}^{b}=\frac{\left|A\right|^{2}}{\sigma^{2}}\sum_{p}^{{N_{{\text{R}}}}}\sum_{m}^{{M_{b}}}\sum_{n}^{{N_{b}}}\left(2\pi mf_{{\text{C}}}^{b}T^{b}\right)^{2}=\frac{4\pi^{2}\left|A\right|^{2}\left(f_{{\text{C}}}^{b}T^{b}\right)^{2}}{\sigma^{2}}\frac{\left(2M_{b}-1\right)\left(M_{b}-1\right)M_{b}N_{{\text{R}}}N_{b}}{6},$ (54)
@@
$F_{\sin\theta_{i,\text{Rx}},\tau_{i,0}}^{b}=F_{\tau_{i,0},\sin\theta_{i,\text{Rx}}}^{b}=\frac{-|A|^{2}(2\pi)^{2}}{\sigma^{2}}\sum_{p}^{N_{\text{R}}}\sum_{m}^{M_{b}}\sum_{n}^{N_{b}}\left(pn\frac{d_{\text{r}}}{\lambda^{b}}\Delta f^{b}\right)=\frac{-4\pi^{2}|A|^{2}\left(\frac{d_{\text{r}}}{\lambda^{b}}\Delta f^{b}\right)}{\sigma^{2}}\frac{\left(N_{\text{R}}-1\right)N_{\text{R}}\left(N_{b}-1\right)N_{b}M_{b}}{4},$ (55)
@@
$F_{\sin\theta_{i,\text{Rx}},\gamma_{i}}^{b}=F_{\gamma_{i},\sin\theta_{\text{Rx}}}^{b}=\frac{|A|^{2}(2\pi)^{2}}{\sigma^{2}}\sum_{p}^{N_{\text{R}}}\sum_{m}^{M_{b}}\sum_{n}^{N_{b}}\left(pm\frac{d_{\text{r}}}{\lambda^{b}}f_{\text{C}}^{b}T^{b}\right)=\frac{4\pi^{2}|A|^{2}\left(\frac{d_{\text{r}}}{\lambda^{b}}f_{\text{C}}^{b}T^{b}\right)}{\sigma^{2}}\frac{\left(N_{\text{R}}-1\right)N_{\text{R}}\left(M_{b}-1\right)M_{b}N_{b}}{4},$ (56)
$F_{\gamma_{i},\tau_{i,0}}^{b}=F_{\tau_{i,0},\gamma_{i}}^{b}=\frac{-(2\pi)^{2}|A|^{2}}{\sigma^{2}}\sum_{p}^{N_{\text{R}}}\sum_{m}^{M_{b}}\sum_{n}^{N_{b}}\left(mn\Delta f^{b}f_{\text{C}}^{b}T^{b}\right)=\frac{-4\pi^{2}|A|^{2}\left(\Delta f^{b}f_{\text{C}}^{b}T^{b}\right)}{\sigma^{2}}\frac{\left(M_{b}-1\right)M_{b}\left(N_{b}-1\right)N_{b}N_{\text{R}}}{4},$ (57)
Combining the relationships $\theta_{i,\text{R}}=\text{arcsin}(\sin\theta_{i,\text{R}})$, $\gamma_{i}=\frac{2v_{i,0}}{c}$, and $\tau_{i,0}=\frac{2r_{i,0}}{c}$, the CRLBs for angle, range, and radial velocity of the CA-enabled MIMO-OFDM system are expressed in (58), (59), and (60), respectively.
$CRLB_{\text{CA}}\left(\theta_{i,\text{Rx}}\right)=\left[\cos\theta_{i,\text{Rx}}\sum_{b=1}^{\mathcal{B}}\mathbf{F}_{b}\right]_{1,1}^{-1},$ (58)
$CRLB_{\text{CA}}\left(r_{i,0}\right)=\left[\left(\frac{2}{c}\right)^{2}\sum_{b=1}^{\mathcal{B}}\mathbf{F}_{b}\right]_{2,2}^{-1},$ (59)
$CRLB_{\text{CA}}\left(v_{i,0}\right)=\left[\left(\frac{2}{c}\right)^{2}\sum_{b=1}^{\mathcal{B}}\mathbf{F}_{1}\right]_{3,3}^{-1}.$ (60)
Upon observing (58), (59), and (60), the CRLB of angle is related to angle, which is consistent with previous work in [43]. Furthermore, the CRLBs of parameter estimations in CA-enabled MIMO-OFDM system decrease with the increase of the number of CCs, revealing the superiority of the proposed ISAC system. It is important to note that there has still a trade-off between the CRLBs of range and velocity, which can be addressed by combining multi-resource optimization or adaptive techniques [45, 21].
@@ V Simulation Results and Analysis
In this section, the simulation results are provided to verify the high-spend communication and high-accuracy sensing capabilities of the proposed CA-enabled MIMO-OFDM ISAC system. Simulation parameters are detailed in Table II. The simulation results are obtained with 1000 times Monte Carlo simulations. It is worth noting that reasonable resource allocation can maximize the benefits of communication and sensing, which is also a focus of future trend. To ensure that the simulation results are not constrained by resource allocation, the simulation considers different bandwidths and different numbers of OFDM symbols on high and low-frequency bands.
TABLE II: Simulation parameters [8, 46, 47, 30, 33].
@@
$f_{\mathrm{C}}^{1}$ Carrier frequency of low-frequency band 3.5 GHz [30, 33] $f_{\mathrm{C}}^{2}$ Carrier frequency of high-frequency band 28 GHz [30, 33]
$M_{1}$ Number of OFDM symbols on low-frequency band 14 [8] $M_{2}$ Number of OFDM symbols on high-frequency band 28
$N_{1}$ Number of subcarriers on low-frequency band 512 [8] $N_{2}$ Number of subcarriers on high-frequency band 512
$T^{1}$ Total symbol duration on low-frequency band 43.9 $\mathrm{\mu}$s $T^{2}$ Total symbol duration on high-frequency band 5.49 $\mathrm{\mu}$s
$\Delta f^{1}$ Subcarrier spacing on low-frequency band 30 kHz [47] $\Delta f^{2}$ Subcarrier spacing on high-frequency band 240 kHz [47]
$N_{\text{T}}$ Number of transmit antennas of BS 128 $N_{\text{R}}$ Number of receive antennas of BS 128
$N_{\text{U}}$ Number of receive antennas of UE {4, 5, 6} $I$ Number of potential targets 3
$\theta_{i,\text{Rx}}$ AoAs of targets $\{30,40,50\}^{\circ}$ $\theta_{i,\text{Tx}}$ AoDs of targets $\{30,40,50\}^{\circ}$
$r_{i,0}$ Ranges of targets {117,150,170} m $v_{i,0}$ Velocities of targets {13,20,25} m/s
@@ V-A Performance of Communication
According to (38) and (41) derived in Section IV-A, the communication MIs of CA-enabled MIMO-OFDM ISAC system and MIMO-OFDM ISAC system under various SNR and $N_{\text{U}}$ are simulated in Fig. 5.
@@
In Fig. 5, the communication MI of the CA-enabled MIMO-OFDM ISAC signal is larger than that of the MIMO-OFDM ISAC signal under different $N_{\text{U}}$. Simultaneously, the communication MI rises with the number of receive antennas $N_{\text{U}}$ in UE, driven by spatial multiplexing in MIMO systems.
@@ V-B Performance of Sensing
In this section, the CRLBs and the average of root mean square errors (ARMSEs) of range and velocity estimations are simulated to demonstrate the high-accuracy sensing performance of the proposed CA-enabled MIMO-OFDM ISAC system. The ARMSEs are denoted by
$\Theta_{r}=\frac{1}{I}\sum_{i=1}^{I}\sqrt{(\hat{r}_{i,0})^{2}-(r_{i,0})^{2}},$ (61)
$\Theta_{v}=\frac{1}{I}\sum_{i=1}^{I}\sqrt{(\hat{v}_{i,0})^{2}-(v_{i,0})^{2}},$
where $\Theta_{r}$ and $\Theta_{v}$ denote the ARMSEs of range and velocity estimations, respectively.
V-B1 Comparison of CRLB
@@
  • Comparing with the CRLB of low-frequency MIMO-OFDM ISAC signal, the high-frequency MIMO-OFDM ISAC signal is lower, which also exist in the OFDM ISAC signals. The reason is as follows. In the same received slots, the number of OFDM symbols in high-frequency band is more than that in low-frequency band, which brings more velocity information of targets.
@@
V-B2 Analysis of CRLB
We further utilize the closed-form CRLB to investigate the relationship between high and low-frequency bandwidth allocation and system sensing performance. Assuming a total available bandwidth of 76.8 MHz, we derive that the subcarrier spacings for high and low-frequency bands satisfy $N_{1}+4\times N_{2}=2560$. The CRLBs for range and velocity estimations are simulated with $N_{2}$ as the variable, as depicted in Fig. 8.
As illustrated in Fig. 8, the bandwidth allocation is categorized into three cases, where the following conclusions are revealed.
  • In Case 1, as the high-frequency bandwidth increases, the CRLB for velocity estimation decreases, whereas the CRLB for range estimation increases. Thus, a trade-off between range and velocity estimation performance emerges.
  • In Case 2, as the high-frequency bandwidth increases, the CRLB for velocity estimation increases, whereas the CRLB for range estimation first increases and then decreases. Hence, reducing the high-frequency bandwidth optimizes overall sensing performance in Case 2.
  • In Case 3, as the high-frequency bandwidth increases, the CRLBs for both velocity and range estimations decrease. Therefore, the high-frequency bandwidth should be increased to its maximum available limit.
The above conclusions provide guidance for the bandwidth allocation of the proposed CA-enabled MIMO-OFDM ISAC system.
Figure 8: CRLBs for different high and low-frequency bandwidth allocations, and the SNR is set as -20 dB
V-B3 ARMSE
The ARMSEs for range and velocity estimations of the proposed CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are simulated in Fig. 9. There are several settings.
@@ V-B3 ARMSE
    The theoretical ARMSEs of range and velocity estimations are expressed in (62) and (63), respectively [48].
@@
Fig. 9(a) reveals the following phenomena.
  • With the increase of SNR, the ARMSEs of CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are stable, because the accuracy of sensing method is limited by the grid size.
  • Comparing with the ARMSEs of MIMO-OFDM ISAC signals, the ARMSE of CA-enabled MIMO-OFDM ISAC signal is lower, indicating the superior performance of range estimation of the proposed CA-enabled MIMO-OFDM ISAC signal over the MIMO-OFDM ISAC signal.
  • When SNR is smaller than $-11.5$ dB, comparing with the ARMSE of low-frequency MIMO-OFDM ISAC signal, the ARMSE of high-frequency MIMO-OFDM ISAC signal is higher. In the simulation, we employed a typical path loss model and set the received SNR of high-frequency signal 5 dB lower than that of low-frequency signal [49]. Therefore, for comparison under the same received SNR, the curve of high-frequency MIMO-OFDM ISAC signal needs to shift 5 dB to the left.
  • With the increase of SNR, the “cliff effect” is occur in ARMSE value, which is common in signal processing. At low SNR, the sensing information is overwhelmed by noise. When the SNR exceeds a certain threshold, the sensing information stands out. This threshold is also commonly used to characterize the robustness of an algorithm.
As shown in Fig. 9(b), the conclusions are similar to range estimation. It should be noted that the higher number of OFDM symbols in the high-frequency band results in a lower ARMSE. Additionally, the absence of the “cliff effect” is because the threshold does not fall within the SNR range of -20 to -5 dB.
(a) ARMSE for range estimation
(b) ARMSE for velocity estimation
Then, the data-level and symbol-level fusion methods are compared in Fig. 10, where the following phenomena are revealed.
  • In Fig. 10, the ARMSE of symbol-level fusion method is consistently lower than that of the data-level fusion method, indicating the superiority of the proposed symbol-level fusion method over the data-level fusion method.
  • With the increase of SNR, the ARMSEs of data-level and symbol-level fusions tend to be flat. This is because the accuracy of sensing is not only affected by bandwidth, SNR and other resources, but also restricted by the grid sizes of range and velocity searching intervals. Therefore, when the influence of SNR is dominant, the curve will continue to decline with the increase of SNR, while when the influence of grid is dominant, the curve will be flat.
  • Compared with the data-level fusion method, the accuracy of range and velocity estimations under symbol-level fusion method can improve by a maximum of 92.77% and 18.31%, respectively.
(a) ARMSE for range estimation
(b) ARMSE for velocity estimation